For a normal distribution, assume the 'sample' standard deviation is 8. What is the maximal margin of error associated with a 99% confidence interval for the true population mean if n = 25. ?

(Round to two decimal places)



1
Expert's answer
2022-05-10T23:40:45-0400

The critical value for "\alpha = 0.01" and "df = n-1 = 24" is "t_c = z_{1-\alpha/2; n-1} = 2.79694."


"E=t_c\times\dfrac{s}{\sqrt{n}}=2.79694\times\dfrac{8}{\sqrt{25}}=4.48"

The maximal margin of error is "4.48."